5 points) What are the frequency. V. and photon ene wavelength of 337.1 nm? and photon energy. E. of light having a
Answer #1

Given value is

wavelength Invalid Equation = 337.1 nm

we know that

1 nm = 1 \times 10-9 m

hence

wavelength Invalid Equation = 337.1 \times 10-9 m

c = speed of light = 3.0  \times 108 m/s

We know the formula

Frequency = speed of light / wavelength

then

Frequency = 3.0  \times 108 m/s / 337.1 \times 10-9 m = 8.90 \times 1014 s-1 = 8.90 \times 1014 Hz

Frequency = 8.9 \times 1014 Hz

Now

Photon energy

E = hc / Invalid Equation

where h = plancks constant = 6.626 \times 10-34 J.s

So

Energy = 6.626 \times 10-34 J.s \times 3.0  \times 108 m/s / 337.1 \times 10-9 m = 5.90 \times 10-17 J

Photon energy = 5.90 \times 10-17 J